OpenStudy (anonymous):

Find the probability of the following five-card poker hands from a 52-card deck. In poker, aces are either high or low. Two pair (2 cards of one value, 2 of another value)

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Poker_probability

OpenStudy (anonymous):

man you got a lot of these huh?

OpenStudy (anonymous):

haaaa yeaaah .

OpenStudy (anonymous):

knock em out

OpenStudy (anonymous):

huuuh ?

OpenStudy (anonymous):

How many ways can the letters in the word "WISCONSIN" be arranged? A. 38,572 B. 45,360 C. 69,620 D. 362,880

OpenStudy (anonymous):

\[\frac{9!}{2\times 2\times 2}\]

OpenStudy (anonymous):

2 "i" 2 "s" 2 "n"

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=9!%2F8 hope that is one of your answers

OpenStudy (anonymous):

How many different sequences of 4 digits are possible if the first digit must be 3, 4, or 5 and if the sequence may not end in 000? Repetition of digits is allowed. A. 1,512 B. 2,999 C. 2,997 D. 2,000

OpenStudy (anonymous):

and it was (: